For the following questions, i) state clearly the shear force and bending moment equations and ii) draw the shear force and bending moment diagrams. Please answer both10 kN 2 kN/m 5 m 3 m —+-2 m 8 kN/m 4 m -
Answer #1

2 kN/m Salt 21Nm í Etter JOEN E MAOP Armas +3m- Am RD RA1 axt V KB (lo-x) 9.5KN Reactions- EFV-0 = RA+ Rp = 2x5 + 10 - 20 kgX SKNlm ť da im kazm &x- AN X (6-2) 16 KN 12 KN X=1.5ml Reactions- Ely : 0 X Rat Rp = 8X6= 48 KN EMAO » 8x666/= RB X 4 - RB =please ask if any doubt

please rate by thumbs up